OpenStudy (anonymous):

Using the public key N = 1763 and k = 23, encode the following message: WICKED.

OpenStudy (across):

Which algorithm?

OpenStudy (anonymous):

wish I knew, that is literally the question given, that's all I'm given, no idea

OpenStudy (cwrw238):

i'll see if i can dig out the book

OpenStudy (anonymous):

thanks! found a Pure Mathematics book which has a part about secret codes, going through that now

OpenStudy (cwrw238):

this is the RSA code which when used with huge numbers is unbreakable - the banks use it to encrypt info on the internet. you encrypt W by using the ASCII representation of W the example in book is X = 1011000 - you of course need to find W you encrypt according to formula C = M^k (mod N) C = ENCRYPTED code, M = letter to be encrypted, k -n your case =23 and N = 1763

OpenStudy (cwrw238):

here M = 1011000 = 88 in decimal so C = 88^23 (mod 1763)

OpenStudy (cwrw238):

shoot! - how do you get 88^23!! - you'll need a massive computer - perhaps i'm not using the correct algorithm - i doubt they'd give you a problem that involves such huge numbers

OpenStudy (cwrw238):

theres 44 digits in 88^23

OpenStudy (anonymous):

thanks though! method in this book has something to do with successive numbers, really confusing stuff

OpenStudy (cwrw238):

no probs - it certainly looks like the RSA code with the public keys. in my book , N is 187 and k = 7 so the arithmetic was easy - the numbers were not so big. but your numbers would need a computer